OpenStudy (anonymous):

Solve the solution 5x+9=39

OpenStudy (anonymous):

Subtract 9 form both sides

OpenStudy (anonymous):

5x + 9 - 9 = 39 - 9 5x = 30 Divide by 5 on both sides x = 6

OpenStudy (shubhamsrg):

 just find the point of intersection.

OpenStudy (anonymous):

Thank you!

OpenStudy (shubhamsrg):

Guess you didn't say thank you to me, but still, you're welcome! ;)